OpenStudy (anonymous):

Will Give Medal To Best Explanation And Answer! Simplify by rationalizing the denominator 4sqrt6 / sqrt30

OpenStudy (anonymous):

I already worked it out, but i want make sure i got it right

OpenStudy (anonymous):

\[(4\sqrt{6})/(\sqrt{30}) = (4\sqrt{6)})/(\sqrt{5\times6})=(4\sqrt{6})/\sqrt{5}\times \sqrt{6}=4/\sqrt{5}\]

OpenStudy (phi):

it is common practice to not leave radicals in the denominator, so multiply by sqrt(5)/sqrt(5)

OpenStudy (anonymous):

why cant i multiply the denominator by it self and the numerator?

OpenStudy (anonymous):

meaning: why cant i multiply the numerator and denominator by sqrt30?

OpenStudy (phi):

you can . you get \[ \frac{4 \sqrt{180} }{30} \] which can be simplified further

OpenStudy (anonymous):

would it change the answer in any way?

OpenStudy (phi):

if the question is simplify, the way you get to the simplified form does not matter. But you are not done yet.

OpenStudy (anonymous):

well all it asks is to simplify, should i keep going to get a completed answer then?

OpenStudy (phi):

I would factor 180 into 2*90= 2*2*45= 2*2*3*3*5 and pull out a 2 and 3 from the sqrt to get 2*3*sqrt(5) then simplify

OpenStudy (anonymous):

if you multiply by sqrt30/sqrt 30 it will be : \[(4\sqrt{6})/(\sqrt{30})\times(\sqrt{30})/(\sqrt{30})=4\sqrt{6}\times \sqrt{30}/(30)=4\sqrt{5}/5\] which's still like what i did

OpenStudy (anonymous):

@phi so i get 6sqrt5 when simplified. it doesnt say to solve tho, should i still?

OpenStudy (phi):

\[ \sqrt{180} = 6 \sqrt{5} \] your problem is \[ \frac{4\cdot \sqrt{180}}{30} =\frac{4\cdot 6\sqrt{5}}{30} = \frac{4\sqrt{5}}{5}\] just like ikarus said.

OpenStudy (phi):

another way to simplify is if you start with \[ \frac{4 \sqrt{6}}{\sqrt{30}}= 4 \sqrt{\frac{6}{30} } = \frac{4}{\sqrt{5}}= \frac{4\sqrt{5}}{5}\]

OpenStudy (anonymous):

oh ok, i think thats what ikarus said originally. Thanks @phi and @ikarus !

OpenStudy (anonymous):

welcome